3. Last year, Jasper Company earned operating income of $29,400 with a contribution margin ratio of 0.3. Actual revenue was $245,000. Calculate the total fixed cost. Note: Round your answer to the nearest dollar, if required..
Answer #1

Contribution margin ratio=Contribution margin/Sales

Hence Contribution margin=(245,000*0.3)=$73500

Contribution margin-Fixed costs=Target operating income

Hence Fixed costs=(73500-29400)

=$44100.